OpenStudy (anonymous):

Sarah has a total of 52 DVD's and CD's. If the number of CD's is two more than four times the number of DVD's, how many CD's does she have?

OpenStudy (anonymous):

CD=42.

OpenStudy (anonymous):

DVD number=10 So 4x10=40, then add 2 more=42. 42+10=52.

OpenStudy (anonymous):

let the number of CD's be x and DVD's be y x=4y+2 x+y=52 now solve

OpenStudy (anonymous):

from the info given we can create a system of equation: "Sarah has a total of 52 DVD's and CD's"=c + d = 52 "If the number of CD's is two more than four times the number of DVD's"=c + d = 52 now solve this system: c + d = 52 c = 4d + 2